$$\ce{2C_4H_{10}(g) +13O_2(g) -> 8 CO_2(g) + 10H_2O(l)}$$
$65cm^3$ of butane is burnt in $65cm^3$ of oxygen. Find the volume of carbon dioxide produced.\

I would like someone to explain this to me.

1
Expert's answer
2022-03-03T04:05:04-0500

If 65cm3 of butane is burnt, number of moles is ;

Moles = mass / rfm

Moles= 65/58= 1.12moles

Moles of carbon produced;

Mole ratio= 1:4

4/5× 1.12= 0.896mole

Mass produced = 0.896× 44= 39.42g

Volume= 39.42cm3